# Repeating nuclear transients from repeating partial tidal disruption events: reproducing ASASSN-14ko and AT2020vdq

### v1

@@ 4.4 Connecting ASASSN-14ko’s outbursts to orbital and stellar properties
For observed events, the recurrence timescale is generally defined as the time between successive peaks in the lightcurve. This timescale does not directly correlate with the orbital or stellar parameters. Our ability to constrain these parameters depends on the presence or absence of certain features in the observed lightcurve. Specifically, there will be some time-lapse between the sequence of events that constitute the process of disruption and the ensuing accretion onto the SMBH, namely, the star reaching pericenter, the initiation of accretion, the rising phase of the lightcurve, and its subsequent decay. If the timescale on which the fallback rate rises and decays is comparable to or longer than the time between successive peaks, then the returning of the stellar core to the pericenter would be associated with a prompt shutoff of the fallback rate (Liu et al., 2023; Wevers et al., 2023), allowing us to constrain the orbital period of the star. Such a prompt shutoff was observed in the X-ray lightcurves of transients such as AT2018fyk and eRASSt-J045650 (Wevers et al., 2023; Liu et al., 2024). On the other hand if, as is the case for ASASSN-14ko, the lightcurve decays on a timescale that is much shorter than the time between successive peaks, then the orbital period of the star can not be constrained from the observed lightcurves of the events. In addition to orbital parameters, rpTDE lightcurves can also allow us to constrain the properties of the disrupted star. As we showed here, ASASSN-14ko’s outbursts are consistent with the partial disruption of a high-mass star with a centrally concentrated core and a diffuse outer envelope, on a grazing orbit around a SMBH, losing $\sim 0.01M_{\odot}$ on each pericenter passages, and surviving multiple encounters to give rise to the observed outbursts.
@@ 5 Summary
We used analytical estimates based on a simplified toy-model, and hydrodynamical simulations of the repeated partial tidal disruption of stars by an SMBH, to analyze the feasibility of the rpTDE model for generating periodic nuclear transients. The key findings of our work are:
@@
    Using the energy-period relation for the Keplerian orbit, we showed that the rotational velocity imparted to the surviving stellar core can lead to changes in its period that are comparable to the observed period derivative of $\dot{P}=-0.0016$ for ASASSN-14ko. Once the star is spun up to $\sim\Omega_{\rm p}$, if additional energy is imparted into oscillatory and quadrupolar oscillations that can be subsequently radiated, this effect would continue on a per-orbit basis.
@@
    The timescale on which the fallback rate of disrupted debris onto the SMBH rises and peaks decreases as a power-law in the imparted angular velocity ($t_{\rm peak}\propto\left(1+\sqrt{2}\Omega/\Omega_{\rm p}\right)^{-0.8}$). It shows greater variation for earlier orbits, and eventually converges to an almost constant value. Thus, if an rpTDE is observed during the first few pericenter passages of the star on its orbit around an SMBH, the timescale on which the lightcurve rises and peaks should show a significant change between successive flares.
@@
We thank Itai Linial and Eliot Quataert for useful discussions. A.B. and E.R.C. acknowledge support from NASA through the Neil Gehrels Swift Guest Investigator Program, proposal number 1922148. E.R.C. acknowledges support from the National Science Foundation through grant AST-2006684, and from NASA through the Astrophysics Theory Program, grant 80NSSC24K0897. C.J.N. acknowledges support from the Science and Technology Facilities Council (grant No. ST/Y000544/1) and from the Leverhulme Trust (grant No. RPG-2021-380). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).

### v2

@@ 4.4 Connecting ASASSN-14ko’s outbursts to orbital and stellar properties
For observed events, the recurrence timescale is generally defined as the time between successive peaks in the lightcurve. This timescale does not directly correlate with the orbital or stellar parameters. Our ability to constrain these parameters depends on the presence or absence of certain features in the observed lightcurve. Specifically, there will be some time-lapse between the sequence of events that constitute the process of disruption and the ensuing accretion onto the SMBH, namely, the star reaching pericenter, the initiation of accretion, the rising phase of the lightcurve, and its subsequent decay. If the timescale on which the fallback rate rises and decays is comparable to or longer than the time between successive peaks, then the returning of the stellar core to the pericenter would be associated with a prompt shutoff of the fallback rate (Liu et al., 2023; Wevers et al., 2023), allowing us to constrain the orbital period of the star. Such a prompt shutoff was observed in the X-ray lightcurves of transients such as AT2018fyk and eRASSt-J045650 (Wevers et al., 2023; Liu et al., 2024b). On the other hand if, as is the case for ASASSN-14ko, the lightcurve decays on a timescale that is much shorter than the time between successive peaks, then the orbital period of the star can not be constrained from the observed lightcurves of the events. In addition to orbital parameters, rpTDE lightcurves can also allow us to constrain the properties of the disrupted star. As we showed here, ASASSN-14ko’s outbursts are consistent with the partial disruption of a high-mass star with a centrally concentrated core and a diffuse outer envelope, on a grazing orbit around a SMBH, losing $\sim 0.01M_{\odot}$ on each pericenter passages, and surviving multiple encounters to give rise to the observed outbursts.
@@ 5 Summary
We used analytical estimates based on a simplified toy model, and hydrodynamical simulations of the repeated partial tidal disruption of stars by an SMBH, to analyze the feasibility of the rpTDE model for generating repeating nuclear transients. The key findings of our work are:
@@
    Using the energy-period relation for the Keplerian orbit, we showed that the rotational velocity imparted to the surviving stellar core can lead to changes in its period that are comparable to the observed period derivative of $\dot{P}=-0.0026$ for ASASSN-14ko. Once the star is spun up to $\sim\Omega_{\rm p}$, if additional energy is imparted into oscillatory and quadrupolar oscillations that can be subsequently lost, this effect would continue on a per-orbit basis. Our arguments (Section 4.3) and simulations (Appendix A) suggest that the mechanism responsible for removing the energy is not predominantly radiative. Instead, it is likely (and consistent with the results in Appendix A) that the thermal energy is deposited in the outer layers of the star that are tidally stripped on subsequent encounters, while the core is relatively unmodified and allowed to survive for mamy more encounters.
@@
    The timescale on which the fallback rate rises and peaks decreases as a power-law in the imparted angular velocity ($t_{\rm peak}\propto\left(1+\sqrt{2}\Omega/\Omega_{\rm p}\right)^{-0.8}$). It shows greater variation for earlier orbits, and eventually converges to an almost constant value. Thus, if an rpTDE is observed during the first few pericenter passages of the star on its orbit around an SMBH, the timescale on which the lightcurve rises and peaks should show a significant change between successive flares.
@@
We thank Itai Linial and Eliot Quataert for useful discussions, and the anonymous referee for useful comments and suggestions that improved the manuscript. A.B. and E.R.C. acknowledge support from NASA through the Neil Gehrels Swift Guest Investigator Program, proposal number 1922148. E.R.C. acknowledges support from the National Science Foundation through grant AST-2006684, and from NASA through the Astrophysics Theory Program, grant 80NSSC24K0897. C.J.N. acknowledges support from the Science and Technology Facilities Council (grant No. ST/Y000544/1) and from the Leverhulme Trust (grant No. RPG-2021-380). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Appendix A Effects of tidal heating
Figure 15: Fallback rates onto a $10^{6}M_{\odot}$ SMBH from the partial disruption of the $3M_{\odot}$ TAMS star on its first six pericenter passages, with (pink) and without shock heating (purple). With the inclusion of shock heating, the outer layers of the star puff-up, leading to an increased amount of mass loss as compared to the non shock heating case. The fallback rates for the second and third encounters show the maximum deviation from the non shock heating case. Once the outer layers of the star are removed, subsequent pericenter passages lead to a comparable amount of mass being lost from the surviving core, and almost indistinguishable fallback rates with and without heating.
To assess the impact of tidal heating on the survivability of the star, we performed simulations for the $3M_{\odot}$ TAMS star that are identical to those presented in Section 3.2, except that the thermal energy that is generated as a byproduct of viscous dissipation (i.e., due to the numerical implementation of an artificial viscosity; see Price et al. 2018 for details in the context of phantom specifically) is retained within the fluid. In contrast, the simulations presented in Section 3.2 assume that this heat is lost from the system; which of these scenarios is closer to reality depends on where the energy is deposited in the star, the rate at which the kinetic energy imparted via tides is viscously and radiatively (or nonlinearly; e.g., McMillan et al. 1987; Kochanek 1992; Kumar & Goodman 1996; Weinberg et al. 2012) damped, and the photon mean free path, but in general we expect the numerical heating to be artificially large in the outer layers of the star where the resolution is lowest (see Norman et al. 2021; Coughlin & Nixon 2022b specifically in the context of TDEs).
The fallback rates for the first six encounters for the 1M particle simulations, with and without “shock heating” (the term we adopt for numerical heating, which is due to the increased importance of viscosity in enforcing the smoothness of the flow, but is not necessarily due to the presence of a shock), are shown Figure 15. While the fallback rates for $N=1$ are almost indistinguishable, the retention of heat within the star causes the outer layers of the star to inflate, and mass is lost more easily from these layers relative to the case where the heat is dissipated. This causes an increased amount of mass loss in the second and third pericenter passages, and the fallback rates for these two encounters lie systematically above the corresponding fallback rates from the simulations in which shock heating is ignored. Table 1 shows the amount of mass lost by the star and that is accreted onto the SMBH for the first six pericenter passages. Once the tidally heated outermost layers of the star are stripped, shock heating is no longer significant and the fallback rates are again indistinguishable.
To test the effect of resolution on our results, we also simulated the first pericenter passage of the star with the inclusion of shock heating at 10M particles. Figure 16 shows the fallback rates with and without shock heating, at the two different resolutions. Aside from differences at the noise level, the curves are identical. We also compared the fallback rates for the $1M_{\odot}$ ZAMS star with $\beta=0.6$ and $\beta=1$, and found no significant differences between the shock heating included and shock heating excluded simulations during the first encounter. Finally, Figure 17 shows the fluid column density projected onto ($x$-$y$) and perpendicular to ($x$-$z$) the orbital plane. While the highest-density regions remain unaltered, the retention of heat causes the outer layers of the star to inflate, which are removed on subsequent encounters.
Figure 16: The fallback rates for the first pericenter passage of the $3M_{\odot}$ TAMS star, with and without shock heating. The solid curves are for a higher resolution of 10M particles, whereas the dashed curves are for a resolution of 1M particles.
Taken together, these tests and the overall insensitivity of our results to the adopted thermodynamic prescription suggest that the star loses the excess heat from tidal dissipation mechanically: it is within the outer layers of the star that most of the thermal energy is deposited, and these layers are subsequently stripped by the black hole. The mechanical (as opposed to radiative) loss of energy is likely required for systems in which the orbital period is as short as ASASSN-14ko’s, because the rate at which energy would need to be radiated – and still yield the observed $\dot{P}$ – would likely be above the Eddington limit of the star. Specifically, from the relationship between the orbital period $T$ and specific energy $\epsilon$ of a Keplerian orbit, we have $\Delta T/T=3\Delta\epsilon/(2\epsilon)$, where $\Delta T$ ($\Delta\epsilon$) is the per-orbit change in the orbital period (specific energy). For $T=114$ days and $M_{\bullet}=10^{7}M_{\odot}$, $\epsilon\simeq 5\times 10^{-4}c^{2}$, and with $\Delta T/T=0.001$, we have $\Delta\epsilon\simeq 8.6\times 10^{-7}c^{2}$. If this energy is placed into the outer $0.01M_{\odot}$ of the star, then the rate at which the energy would need to be lost over the orbital period of the star is $\dot{E}=0.01M_{\odot}\times\Delta\epsilon/T\simeq 1.5\times 10^{39}$ erg s-1 $\simeq 3.5L_{\rm Edd}$, where $L_{\rm Edd}=4\pi GM_{\star}c/\kappa$ is the Eddington luminosity with $M_{\star}=3M_{\odot}$ and $\kappa=\kappa_{\rm es}=0.34$ cm2 g-1 for solar abundances. Since the thermal diffusion timescale is $\sim R_{\star}/(c/\tau)\gg 114$ days, where $\tau$ is the optical depth over the stellar radius, the tidal energy can be deposited and stored in the outer layers of the star (which is also where the amplitude of the tidal force is largest and hence where we expect most of the kinetic energy to be nonlinearly and/or viscously damped), thus allowing the core of the star to survive and giving rise to ASASSN-14o-like flares over a prolonged period.
Figure 17: The column density profile of the $3M_{\odot}$ TAMS star at $\sim 1$ day past pericenter passage, with (right) and without (left) shock heating, at a resolution of 10M particles. The top panel shows a comparison of the density profile in $x-y$ coordinates, and the bottom panel shows the density profile in $x-z$ coordinates. The outer layers of the star are inflated in the simulation including shock heating, but structure of the core is largely unaffected, thus preventing the star from being completely disrupted in a few orbits.
Pericenter Passages Mass Stripped $\left[M_{\odot}\right]$ Mass Accreted $\left[M_{\odot}\right]$
without shock heating    with shock heating without shock heating    with shock heating
1 $1.65\times 10^{-2}$                        $1.74\times 10^{-2}$ $8.46\times 10^{-3}$                        $8.53\times 10^{-3}$
2 $1.47\times 10^{-2}$                        $2.06\times 10^{-2}$ $7.54\times 10^{-3}$                        $1.06\times 10^{-2}$
3 $1.31\times 10^{-2}$                        $1.89\times 10^{-2}$ $6.45\times 10^{-3}$                        $9.44\times 10^{-3}$
4 $1.19\times 10^{-2}$                        $1.43\times 10^{-2}$ $5.95\times 10^{-3}$                        $8.46\times 10^{-3}$
5 $1.12\times 10^{-2}$                        $1.23\times 10^{-2}$ $5.55\times 10^{-3}$                        $6.27\times 10^{-3}$
6 $1.07\times 10^{-2}$                        $1.12\times 10^{-2}$ $5.37\times 10^{-3}$                        $5.71\times 10^{-3}$
Table 1: Amount of mass lost per encounter by the $3M_{\odot}$ TAMS star orbiting a $10^{6}M_{\odot}$ on a $\beta=1$ orbit, and mass accreted onto the SMBH, with and without the inclusion of shock heating.
Appendix B Development of Solid-Body Rotation in the Core
To address the temporal evolution and resolution dependence of the angular velocity of the partially disrupted star, Figure 18 shows the angular velocity of the stellar core as a function of time. Here $t_{\rm pericenter}$ is the time taken for a Keplerian orbit to reach pericenter from the initial location of the stellar COM (equal to $5r_{\rm t}$). The angular velocity $\Omega$ is calculated as an average over every particle present within a radius $R_{\rm c}\sim 0.2R_{\odot}$ from the center of the star, i.e., within the “core” as defined by Coughlin & Nixon (2022a). As seen from the figure, the star is spun up to approximately its asymptotic value on $\sim$ the stellar dynamical time upon reaching pericenter, though there is some temporal evolution on longer timescales that likely arises from the re-accretion (by the core) of a fraction of the tidally stripped tails. This figure also shows the angular velocity from the 10M particle simulation to $\sim 1$ day past pericenter, which is in good agreement with the lower-resolution simulation. The solid-body rotation that is induced in the star is therefore resolved and physical.
Figure 18: Rotational velocity $\Omega$ imparted to the core of the $3M_{\odot}$ TAMS star, in units of its breakup spin, as a function of time measured relative to when the COM of the star reaches pericenter. The initially non-rotating star is spun-up near pericenter, and continues to rotate with a roughly constant angular velocity beyond $\sim 2$ days past pericenter.
The emergence of bulk rotation as a consequence of the excitation of quadrupolar modes was studied by  Kochanek (1992), who pointed out that the quadrupolar $\ell=2,m=-2$ mode is tidally excited with a higher amplitude compared to other oscillatory modes, and the conservation of circulation within the star requires that the damping of this mode (which possesses both angular momentum and vorticity) generates a bulk rotation within the star at an angular frequency that is much smaller than the mode frequency (see the discussion in Section 3 of  Kochanek (1992)). Our simulations show that tidal interactions spin up the star to a uniform angular velocity, which is a small fraction of $\Omega_{\rm p}$. The rotational energy thus imparted to the core does not exceed its binding energy, and also does not account for the dissipation of orbital energy (as we show in Appendix A above, the energy imparted through tides heats up the outer layers of the star, which are mechanically removed on subsequent encounters with the SMBH).